The length and breadth and thickness of rectangular sheet of wood are 3.230m 1.49m and 1.99m respectively determine the surface and volume of the sheet up to correct significant figures


1
Expert's answer
2021-10-06T15:52:36-0400

The volume is given as follows:


"V = 3.230m\cdot 1.49m\cdot 1.99m\approx 9.58m^3"

The surface consists of pairs of three following parts:


"A_1 = 3.230m\cdot 1.49m\\
A_2 = 1.99m\cdot 1.49m\\
A_3 = 3.230m\cdot 1.99m"

Finally:


"A = 2(A_1 + A_2 + A_3)\\
A = 2\cdot (3.230m\cdot 1.49m + 1.99m\cdot 1.49m + 3.230m\cdot 1.99m) \approx 14.21m^2"

Answer. Volume: 9.58m^3, Surface: 14.21m^2.